A gaseous hydrocarbon gives upon combustion 0.72g of water and 3.08g of carbon dioxide.what will be the empirical formula of hydrocarbon?

Grade:12th pass

1 Answers

Parth Chhabra
6 years ago
A hydrocarbon is consist of C and H and empirical formula can be written as CxHy.To find the empirical formula, you have to find the moles of C and H relesed in form of H2O and CO2.
18 g H2O contain 2 moles of H atoms
0.72 gram of H2O will contain = 2/18 x 0.72 = 0.08 moles of H
​44 g of CO2 contain 1 mole of C atom 
3.08 gram of CO2 will contain 1/44 x 3.08 = 0.07 moles of C
​Ratio of x and y = 0.07/0.08 = 7 : 8
So the empirical formula will be: C7H8
HOPE YOU GET IT